OpenStudy (anonymous):

What is the approximate circumference of a circle with a radius of 40 inches? Use π ≈ 3.14. A. 125.6 inches B. 251.2 inches C. 502.4 inches D. 5024 inches

OpenStudy (anonymous):

circumference = 2*pi*r

OpenStudy (calculusxy):

Simply: pi*diameter

OpenStudy (anonymous):

= 2*3.14*40 =6.28*40 = 251.20

OpenStudy (anonymous):

We need to use this formula C= 2* pi *r Here r = 40 plug this value into furmula

OpenStudy (calculusxy):

You just multiply the radius by 2 to get the diameter.

OpenStudy (anonymous):

ok thank you all

OpenStudy (calculusxy):

You're welcome